OpenStudy (anonymous):

Help graphing linear inequality 5x-y<=6

jimthompson5910 (jim_thompson5910):

5x-y<=6 -y <= -5x + 6 y >= 5x - 6

jimthompson5910 (jim_thompson5910):

So 5x-y<=6 is the same as y >= 5x - 6

jimthompson5910 (jim_thompson5910):

To graph y >= 5x - 6, you graph the line y = 5x - 6 Make it a solid line. Then you shade above the line (since we have a >= sign)

OpenStudy (anonymous):

Thanks. That made a lot of sense.

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

What about one that looks like x<=-2y

jimthompson5910 (jim_thompson5910):

Divide both sides by -2 (the sign will flip) to isolate y x<=-2y x/(-2) >= y y <= -1/2x So graph y = -1/2x, make it a solid line. Then shade below the line.